OpenStudy (anonymous):

if f(x)=x^2-3x and g(x)=√x-1 then f(g(x))

OpenStudy (anonymous):

You'd have\[(\sqrt{x}-1)^{2}-3x\]I'm not sure it's its sqrt x or sqrt x-1, so you may have to change that.

OpenStudy (australopithecus):

f(x)=x^2-3x and g(x)=√x-1 so, f(g(x)) = ((x-1)^(1/2))^2 - 3(x-1)^(1/2) or \[f(g(x)) = (\sqrt{x-1})^2 - 3\sqrt{x-1}\]

OpenStudy (australopithecus):

why did you ask the same question over again?

OpenStudy (anonymous):

what comes after that?

OpenStudy (australopithecus):

remember, Rules: \[\sqrt{x} =x^{\frac{1}{2}}\] \[x^{a}x^{b} = x^{a + b}\]

OpenStudy (australopithecus):

for example, \[x^{2}x^{3} = x^{2+3} = x^{5}\] Also note that, \[x^{1} = x\] And just so you understand \[x^2 = x*x\] \[x^3 = x*x*x\] etc.

OpenStudy (anonymous):

oh i forgot about that, but i have remenber now, but i still get the worng answer

OpenStudy (australopithecus):

so first write this as it is its expanded form based on the definition of an exponent (last rule I showed you) \[(\sqrt{x-1})^2 = ((x-1)^{\frac{1}{2}})^2\]

OpenStudy (australopithecus):

they expect simplification right?

OpenStudy (australopithecus):

I hope you understand what I mean? If not I can just write it out for you

OpenStudy (anonymous):

please

OpenStudy (australopithecus):

Fine fine, \[((x-1)^{\frac{1}{2}})^2 = (x-1)^{\frac{1}{2}}*(x-1)^{\frac{1}{2}}\]

OpenStudy (australopithecus):

do you see what I'm doing? and how to simplify this?

OpenStudy (australopithecus):

all the rules I gave you apply to your problem

OpenStudy (anonymous):

wait is √x then -1

OpenStudy (anonymous):

is not together

OpenStudy (australopithecus):

is it? \[\sqrt{x} - 1\] or is it? \[\sqrt{x - 1}\]

OpenStudy (australopithecus):

sorry but in my defence it was ambiguous, next time make sure you use brackets

OpenStudy (anonymous):

the first one, sorry

OpenStudy (australopithecus):

so can you write the expression out? I can go over the simplification with you

OpenStudy (australopithecus):

you know what f(g(x)) is right?

OpenStudy (australopithecus):

you just replace x with (x)^(1/2) - 1 in f(x)

OpenStudy (anonymous):

yes i know all of that ill write it out over again

OpenStudy (anonymous):

if f(x)=x^2-3x and g(x)=√x -1 then f(g(x))

OpenStudy (australopithecus):

so, \[(x^\frac{1}{2} - 1)^2 - 3(x^\frac{1}{2} - 1)\]

OpenStudy (anonymous):

ok

OpenStudy (australopithecus):

they want you to simplify it right?

OpenStudy (australopithecus):

so first expand the exponent on the first term in f(g(x)) like I did previously

OpenStudy (anonymous):

the answer is x-5√x +4 , but i dont know how to get there

OpenStudy (australopithecus):

They want you to simplify it so first expand the exponent in f(g(x))

OpenStudy (australopithecus):

remember the rule, x^2 = x*x x^3 = x*x*x etc... similarly (x+1)^2 = (x+1)(x+1) (x+1)^3 = (x+1)(x+1)(x+1) etc...

OpenStudy (australopithecus):

so first off you want to expand all your whole number exponents

OpenStudy (australopithecus):

look back to when I thought g(x) = (x -1)^(1/2)

OpenStudy (anonymous):

can you do it please, im getting confused

OpenStudy (australopithecus):

I'm not here to do your work for you, I will only guide you. Math is confusing but just be patient and you will understand it. sorry im not being clear I will try to be more clear from here on out. first step expand this term in the equation ((x)^(1/2) - 1)^2 this is an example of an expansion: (x+1)^3 = (x+1)(x+1)(x+1)

OpenStudy (australopithecus):

I want you to learn this stuff and understand it so you dont struggle with it because it isnt going away. Algebra is actually very simple once you understand the rules

OpenStudy (australopithecus):

can you show me the first step?

OpenStudy (australopithecus):

\[((x)^{\frac{1}{2}} - 1)^2 = ?\] just expand this that is your first step

OpenStudy (anonymous):

i know math is just this equation that i dont remenber, i dont have your time, i skiped this one to do the rest of my equations

OpenStudy (australopithecus):

algebra is just a few simple rules, just please try expanding the term I showed you I will help you with the next step if you do, just make a guess

OpenStudy (anonymous):

nevermine ,

OpenStudy (australopithecus):

there is no shame in being wrong, please dont give up

OpenStudy (australopithecus):

If you get it wrong I will give you the right answer

OpenStudy (australopithecus):

Just work with me try to apply the rules I show you

OpenStudy (anonymous):

hey im not giving up i already got it when you told me the about the square root than converts to on half , you make this too long

OpenStudy (anonymous):

dont anwer

OpenStudy (australopithecus):

Im not making this long I'm only asking you to go through the simplification step by step or you could show me your work

OpenStudy (australopithecus):

instead you just argue with me.

OpenStudy (australopithecus):

what ever I dont have time for this here is the answer

OpenStudy (australopithecus):

First step expand: \[(x^{\frac{1}{2}} -1)^2 -3(x^{\frac{1}{2}} - 1) = (x^{\frac{1}{2}} - 1)(x^{\frac{1}{2}} - 1) - 3(x^{\frac{1}{2}} - 1) \] Second step you distribute (a + b)(d + c) = ad + ac + bd + bc \[(x^{\frac{1}{2}} - 1)(x^{\frac{1}{2}} - 1) - 3(x^{\frac{1}{2}} - 1) = x^{\frac{1}{2}}x^{\frac{1}{2}} - x^{\frac{1}{2}} - x^{\frac{1}{2}} + (-1)(-1) - 3(x^{\frac{1}{2}} - 1)\] Third step distribute the 3 so b(a+c) = ba + bc \[x^{\frac{1}{2}}x^{\frac{1}{2}} - x^{\frac{1}{2}} - x^{\frac{1}{2}} + 1 + (- 3)x^{\frac{1}{2}} + (- 1)(-3)\] forth step collect terms also add exppnents \[x^{\frac{1}{2}}x^{\frac{1}{2}} - x^{\frac{1}{2}} - x^{\frac{1}{2}} + 1 - 3x^{\frac{1}{2}} + 3 = x^{\frac{1}{2} + \frac{1}{2}} - 2x^{\frac{1}{2}} - 3x^{\frac{1}{2}} + 4\] \[x^{1} - 5x^{\frac{1}{2}} + 4 = x- 5x^{\frac{1}{2}} + 4\]

OpenStudy (australopithecus):

Hope this helps probably how I should have went about it in the first place

OpenStudy (australopithecus):

You probably wont even seen it but whatevers